OpenStudy (anonymous):

1 and two-fifths times 4 and two-thirds equals 4 and four-fifteenths 4 and one-half 6 and eight-fifteenths eleven and one-fifth

OpenStudy (anonymous):

Okay so you start with turning them into improper fractions(1numerator and 1 denominator) 1 2/5=7/5, 4 2/3=14/3 7/5 *14/3 =Product of numerators/product of numerators =14*7 / 5*3 = 98*/15 Do the same with the rest Hope I answered your question

OpenStudy (anonymous):

add the wut multiply wut fraction who?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You turn 1 2/5 and 4 2/3 to an improper fraction

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

As an improper fraction they are 7/5 and 14/3

OpenStudy (anonymous):

7/5 14/3

OpenStudy (anonymous):

/yea

OpenStudy (anonymous):

Yes ok then you times that so 7/5 times 14/3 =

OpenStudy (anonymous):

and, the lcd would be 15?

OpenStudy (anonymous):

wut?

OpenStudy (anonymous):

Yah it would be 15

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Now whats 7 times 14?

OpenStudy (anonymous):

umm, where did 14 come from?

OpenStudy (anonymous):

nver mind

OpenStudy (anonymous):

WE made it improper so what is 14 times 7 zambie?

OpenStudy (anonymous):

98

OpenStudy (anonymous):

Yes so you got your answer the answer is 98 / 15

OpenStudy (anonymous):

:D thnx

OpenStudy (anonymous):

medal 4 u :D

OpenStudy (anonymous):

You welcome! You did awesome sauce!

OpenStudy (anonymous):

u got a new fan

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

AWWW THANK YOU byes

OpenStudy (anonymous):

:b